Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: B configuration move ongoing. Not strong science need during the move so ideal time for testing efforts. Some stability loss recently with hardware/software updates.
  • Call for proposals (Claire): 193 EVLA proposals! (normally 120-130) - first for semester system. March 22nd for technical reviews. End of May for notification.
  • ApJ Special EVLA Issue: 53 responses; there will be a note in the e-news that extends interest to 15 February; estimates are that we'll get 75% of those articles. Instructions/timeline to come (expect mid-April deadline for publishing in June/July).
  • Key dates:
    • 09-23 Feb: 4-band test installation
    • 14-18 Feb: Meerkat visitors (Operations)
    • 16 Feb: Correlator Fire Suppression tests
    • 07-10 Mar: New Worlds, New Horizons (Santa Fe)
    • 07 May - 15 Jun: 4-band science observing
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC
    • 22 Feb - 08 Mar: Sonja Vrcic (Correlator)
    • 28 Feb - 28 Mar: Bob Sault

Report from Dish Verification Meeting

  • DVA1 - Dish Verification Antenna 1 (SKA TDP=technology development program) - work toward a Pathfinder SKA prototype; Conceptual design review - get recommendation on whether to proceed or kill it (i.e., don't build an antenna). Napier, Woody, DeBoer, Goldsmith - recommended: significant technical design progress to warrant proceeding. NRAO leadership needed during construction phase of this project; needs an agreement between Fred/Jim Cordes (Cornell); Bob will write MOU. Rick notes: they are looking for site for construction for this summer (20 GHz top frequency - 15 m offset (15x18m)). Barn is not big enough. Needs a building to construct the dish.

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon-Tue, Thu-Fri:
    • 0900-1700 Testing/software development time
    • 1800-0900: Science operations
  • Wed: Maintenance day

Issues for Science Observing:
  • Correlator: fshift sub-band anti-aliasing causing phantom spectral lines when using recirculation for narrow bands (only worry if Michael has to prep your file)
  • Correlator: Missing/mislabeled scans in data sets (small percentage but recurring).
  • Correlator: time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags.
  • Construction: RFI sweep at Ku, K, Ka, Q pre-requisite for high pass filters for use with L301.
  • Calibration: Switched power heuristics for reduction.
  • Array/Construction: Dhawan/Rupen: RF switch issues; new set coming to be tested.
  • Correlator: Delay rocking across bands.
  • Correlator: Subset of BlBs misbehaving (only affects hardware stacking modes)
  • Array: Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • *Correlator*: missing sub-bands (initially thought tied to the station board upgrades; backed out); still occurring through 29 Jan.

Hardware status and updates (Langley, Perley, Rupen)

Software status and updates (Butler)

  • Links:
  • OPT webtest updated (ability to export again); Executor update.
  • CASA Update (new functionality available/issues/enhancements):Notes (Miriam): casapy-test is fixed and recommended (fix affected polarization angles). FO noted that he is still having some issues with UVFITS. If have issues with missing scans see CAS-2779 for information.
  • AIPS Update (new functionality available/issues/enhancements): Notes: Eric is working more on comparing weights; currently TYAPL and REWAY generate different weights (different by an order of magnitude). Behaviors are self consistent.
  • SDM Update: Notes:Continuing work to establish paths for CASA/AIPS use of switched power/cal device tables.

Commissioning Reports

High/Mid Frequency: Canon: Ku-band reference pointing tests

High Frequency/Correlator: Dhawan: Time-averaging CX OSRO/RSRO reference pointing tests

Group updates (Algorithm, Low Frequency, Polarization)

Commissioning Group updates and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

RSRO/ECSO Reports

Individual reports (updates)

  • Leo Ring update (AW775); Brian Kent and Aeree Chung are working on the data. Exciting early images.
  • Brown dwarf pulsations; Greg presented results stored at: https://bugs.nrao.edu/browse/RSRO-27

Misc:

Demo Science Status

AOB

  • Shift log tool

Misc Actions

  • [Frail/Butler]: Enable restrictions based on proximity to sun; review OST weather trending.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 2011-02-07
Topic revision: r5 - 2011-02-08, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback